# VGC: A High-Performance Zone-Based Garbage Collector Architecture for Python with Partitioning and Parallel Execution

### v1

@@ Abstract
The Virtual Garbage Collector (VGC) introduces a novel memory management framework designed to optimize performance across diverse systems, ranging from resource-constrained embedded devices to high-performance parallel architectures. Unlike conventional garbage collectors, VGC employs a dual-layer architecture—Active VGC and Passive VGC—to enable efficient, low-overhead memory management. Active VGC dynamically manages runtime objects using a concurrent mark-and-sweep strategy tailored for parallel workloads, reducing pause times by up to 30% compared to generational collectors in multithreaded benchmarks. Passive VGC, which operates at compile time, optimizes static object allocation through predictive memory mapping, minimizing fragmentation by aligning objects to cache boundaries. This separation of responsibilities ensures predictable memory access patterns, reduces total memory usage by up to 25%, and improves scalability for modern parallel applications. By integrating compile-time and run-time optimizations, VGC offers a robust, adaptable solution for memory-intensive systems, achieving superior performance and reliability in both low-level and high-level programming environments.
@@ 1 Introduction
Garbage collection is a fundamental component of modern programming languages, automating memory management to prevent resource leaks and enhance software reliability. In Python, however, the garbage collector’s reliance on reference counting and the Global Interpreter Lock (GIL)[3], introduces several performance bottlenecks, including increased memory overhead, excessive fragmentation, unpredictable collection pauses, and limited support for true parallel execution. These constraints significantly reduce Python’s scalability in high-performance computing scenarios, particularly in data-intensive and multithreaded workloads.Classical garbage collection techniques rely on heap traversal, reference counting, or generational heuristics, which introduce non-deterministic pause behavior and scaling limitations [5], To address these challenges, we propose the Virtual Garbage Collector (VGC), a novel memory management framework designed to augment or serve as an alternative to Python’s GIL-based garbage collector. VGC employs a dual-layer architecture consisting of Active VGC, which manages dynamic objects at runtime, and Passive VGC, which optimizes the allocation of static objects at compile time. This separation enables fine-grained and efficient memory handling by classifying objects into three distinct zones—Red, Green, and Blue—based on usage frequency and computational complexity. Frequently accessed objects are prioritized in the Green Zone, while rarely used or high-complexity objects are allocated to the Blue or Red Zones, reducing memory fragmentation and improving cache locality. VGC introduces a checkpoint-based bitfield tracking system that replaces per-object reference counting, allowing object states to be evaluated using bitwise logic operations (e.g., AND, OR, NOT) in constant time. This eliminates recursive traversal during garbage collection and reduces latency by up to 20% in initial benchmarks. Furthermore, a lightweight Yield Memory layer handles small-scale operations, such as primitive variable assignments and basic arithmetic, without triggering full garbage collection cycles to further minimize overhead. Unlike conventional garbage collectors (e.g., Java’s CMS or Go’s concurrent GC), VGC’s triadic zoning and checkpoint-based design provide a predictable, low-latency, and scalable memory management solution suitable for parallel workloads. This paper presents the architecture and operational principles of VGC, its RGB zoning strategy, and benchmark results demonstrating its advantages over Python’s GIL-based garbage collector. By integrating partitioned memory zones, bitfield tracking, and yield-based optimization, VGC achieves reduced memory usage, lower fragmentation, and enhanced scalability, paving the way for efficient memory management in modern computing environments.
@@ 2 System Architecture
The architecture of the Virtual Garbage Collector (VGC) is designed around a layered, zone-oriented framework that emphasizes efficiency, determinism, and parallel scalability. Unlike traditional garbage collectors, which depend on recursive traversal and per-object reference counting, VGC employs a compact bit addressing and bitfield checkpoint system that allows constant-time object state evaluation. This design enables VGC to act both as an alternative and as a parallel companion subsystem to the default Python memory manager.
@@ 2.1 Overview of Architectural Layers
VGC consists of four key layers, each responsible for a distinct phase of memory lifecycle management:
@@
    Memory Zoning Layer (R/G/B): All allocated objects are stored in one of three predefined memory zones—Red, Green, or Blue. This classification is based on access frequency, object complexity, and lifetime. Each zone functions as a semi-independent allocator with its own internal management structures, minimizing interference between workloads of different types.
@@
    Checkpoint Bitfield Layer: Instead of storing reference counts in each object header, VGC maintains a centralized checkpoint table in which each object is represented by a single bit or a small fixed bit pattern. These bits encode object states such as active, idle, recyclable, or scheduled for collection. State evaluation and updates are performed through logic gate operations (AND, OR, NOT, XOR, XNOR, NOR, NAND), enabling constant-time liveness detection.
@@
    Yield Memory Layer: This lightweight execution layer acts as a short-term buffer or workspace for small-scale operations such as variable assignments, simple arithmetic, and temporary objects. Yield memory can operate independently of the checkpoint and zoning layers for ephemeral workloads, reducing unnecessary GC overhead. Once execution completes, yield memory either discards the object or notifies the checkpoint system to register persistent values.
@@
    Control and Execution Layer: This layer orchestrates the interaction between zones, checkpoint logic, and yield memory. It includes scheduling logic, partitioning control for multi-core workloads, and optional active-passive GC layering. Active VGC manages runtime and dynamic workloads, whereas Passive VGC optimizes static allocations at compile time.
@@ 3.2 Yield Memory Execution Path
For ephemeral operations such as:
x = 85 // a random variable
the standard garbage collection pipeline is bypassed. Yield Memory allocates a minimal execution unit, executes the operation, and immediately releases the allocation.
By skipping both checkpoint evaluation and VGC zone allocation for short-lived workloads, Yield Memory provides a fast execution path that reduces allocation overhead and runtime latency.
If persistence is required, Yield Memory notifies the checkpoint system, which assigns a 3-bit liveness state and registers the object into the appropriate VGC zone (typically Green).
@@ 3.3 Three-Bit Checkpoint Architecture
The checkpoint layer forms the core of VGC’s memory tracking system. Unlike conventional reference counting, which maintains an integer per object, VGC uses a compact 3-bit encoding for each object’s state. This allows for richer state representation while preserving constant-time evaluation. This 3-bit architecture allows VGC to precisely track object lifecycle transitions without per-object counters. State changes are performed via bitwise logic operations rather than recursive reference updates.
@@ 3.4 Object Feature Profiling and Runtime Metrics
  • $\rho(o)$Referential Fan-Out: The number of distinct inbound references point to the object, representing its connectivity within the program’s object graph.
@@ 3.10 Parallelism and Partitioning
VGC integrates a partitioning model that distributes workloads across CPU cores, with zones operating independently. Checkpoint bitfield updates are parallelized, avoiding locking overhead typical in Python’s GIL. Because the checkpoint system is based on 3-bit state logic, partitioning can use bitmask operations to batch-evaluate objects across zones concurrently. This allows multiple threads or interpreters to process memory in parallel without race conditions or serialized refcount updates. In summary, VGC’s working principle combines three-bit checkpoint state encoding, zone-based allocation, logic-gate-driven decision making, lambda-optimized yield execution, and parallel Execution by partitioning workloads. This results in a deterministic, low-latency, and highly scalable alternative to Python’s GIL-based garbage collector.
@@ 4 Lambda and Three-Bit Checkpoint Mapping
5 Logic Gate–Based Architecture
The Logic Gate–Based Architecture[5], forms the computational core of the VGC checkpoint system. Rather than relying on iterative reference tracking or recursive object scanning, VGC encodes each object’s state within a compact 3-bit field, which is processed directly through gate-level logic operations. This design enables constant-time evaluation of object lifecycles and eliminates the overhead associated with conditional branching or traversal-based garbage collection.
@@ 5.1 Checkpoint Logic Integration
Each allocated object is represented by a 3-bit checkpoint code, denoted as:
@@
where each bit corresponds to a specific state flag used to determine object liveness and collection readiness. The checkpoint bitfields are grouped into logical blocks, allowing simultaneous evaluation of multiple objects using bitwise operations. Instead of maintaining numerical reference counts, the system interprets combinations of bits through logic gates that perform rapid evaluation at the binary level.
  • The AND gate ($A\land B$) ensures that an object remains valid only if both the current state and zone permission bits are active. It acts as a confirmation gate for “keep-alive” conditions.
  • The OR gate ($A\lor B$) preserves objects that are still linked to high-priority zones (such as Green or Blue) even when their primary reference becomes inactive.
  • The NOT gate ($\lnot A$) is used to invert inactive bits, identifying idle or expired objects ready for reclamation.
  • The XOR gate ($A\oplus B$) tracks state transitions by detecting temporal differences between consecutive epochs of the same object.
  • The XNOR gate ($\overline{A\oplus B}$) identifies stability across multiple evaluation cycles, ensuring that persistent objects are not prematurely collected.
  • The NAND and NOR gates are used in composite checkpoint circuits, enabling conditional sweeping and deferred reclamation for objects marked with state patterns $101$ (Deferred) or $110$ (Marked).
@@ 6.2 Core Partitioning Principle
where $P$ denotes the partition ratio representing the load distribution per core. If $T>C$, VGC dynamically clusters multiple Threads under a single core using an event-driven scheduler; if $T\leq C$, interpreter are mapped in a one-to-one configuration to maximize concurrency.
@@ 6.3 Zone-Based Partition Allocation
where $p$ represents a partition index. Each partition execute independently under a dedicated thread or lightweight core instance. The checkpoint and yield layers synchronize periodically to maintain zone integrity without introducing global locking mechanisms.
@@ 6.4 Parallel Execution Interpreter Synchronization
where $C_{T}$ denotes the computed checkpoint state for Thread $T$. Once local evaluations complete, the results are merged via a Sub-Allocator Aggregator that reorders and combines results from multiple threads and core into a coherent memory snapshot.
This mechanism replaces Python’s traditional Global Interpreter Lock (GIL) with a non-blocking, partition-aware synchronization system capable of true multi-core concurrency.
@@ 6.6 Multi-Layer Parallel Model
  • Inter-Zone Parallelism: Red, Green, and Blue zones execute concurrently under separate threads or core, merging results through the checkpoint synchronization layer.
@@ 8.7 Single core : No Partition 40K logical steps (deep recursion)
Table 2.7: Single-core PPE Deep Recursion Benchmark (40000 logical steps, recursion depth per chunk: 4000)
@@ 8.8 Single core : No Partition 80K logical steps (deep recursion)
Table 2.8: Single-core PPE Deep Recursion Benchmark (80000 logical steps, recursion depth per chunk: 8000)
@@ 8.9 Single core : No Partition 160K logical steps (deep recursion)
Table 2.9: Single-core PPE Deep Recursion Benchmark (160000 logical steps, recursion depth per chunk: 16000)
@@ 8.19 Dual core : 2 Partitions 100K logical steps (deep recursion)
Table 2.19: Dual-core PPE Deep Recursion Benchmark (100000 logical steps, recursion depth per chunk: 10000)
@@ 8.20 Dual core : 2 Partitions 200K logical steps (deep recursion)
Table 2.20: Dual-core PPE Deep Recursion Benchmark (200000 logical steps, recursion depth per chunk: 20000)
@@ 8.21 Dual core : 2 Partitions 400K logical steps (deep recursion)
Table 2.21: Dual-core PPE Deep Recursion Benchmark (400000 logical steps, recursion depth per chunk: 40000)

### v2

# Virtual Garbage Collector (VGC): A Zone-Based Garbage Collection Architecture for Python’s Parallel Runtime

@@ Abstract
The Virtual Garbage Collector (VGC) proposes a zone-based memory management architecture aimed at improving execution predictability and memory behavior in Python runtimes. The design explores a dual-layer model consisting of an Active VGC, responsible for managing runtime object lifecycles, and a Passive VGC, intended as a compile-time optimization layer for static allocation planning.
Rather than relying on traditional heap traversal or generational heuristics, VGC introduces memory zoning and checkpoint-based state evaluation to reduce allocation churn and constrain garbage collection scope. Execution partitioning is experimentally evaluated to isolate workloads and localize memory pressure, enabling more deterministic behavior under loop-intensive, recursive, and compute-heavy workloads.
This work presents the architectural principles, execution model, and experimental observations of VGC within a partition-aware runtime context. While the full realization of the dual-layer design is an ongoing effort, the results indicate that zone-based allocation and partitioned execution provide a viable foundation for improving scalability and memory predictability in Python-oriented systems.
@@ 1 Introduction
Garbage collection is a fundamental component of modern programming languages, automating memory management to prevent resource leaks and improve software reliability. In Python, memory management relies heavily on reference counting and the Global Interpreter Lock (GIL)[3], which together introduce several well-known limitations, including increased memory overhead, fragmentation, non-deterministic collection pauses, and restricted scalability for CPU-bound workloads. These constraints become more pronounced in data-intensive, loop-heavy, and multithreaded execution scenarios.
Classical garbage collection techniques typically rely on heap traversal, reference counting, or generational heuristics, all of which incur runtime overhead and introduce variability in pause behavior [5]. While such approaches are effective for general-purpose workloads, they provide limited control over object lifecycles and memory locality, particularly in high-performance and parallel execution contexts.
To explore alternative design choices, this paper proposes the Virtual Garbage Collector (VGC), a zone-based memory management framework designed to improve execution predictability and memory behavior in Python-oriented runtimes. VGC introduces a dual-layer architectural model consisting of an Active VGC, responsible for managing runtime object lifecycles, and a Passive VGC, intended as a compile-time mechanism for guiding static allocation decisions. The separation of responsibilities allows memory behavior to be reasoned about more explicitly across different execution phases.
At the core of VGC is a memory zoning strategy that classifies objects into three predefined zones—Red, Green, and Blue—based on observed access patterns and expected computational characteristics. Frequently accessed objects are assigned to the Green zone, while less frequently accessed or higher-cost objects are placed in the Red or Blue zones. This zoning model aims to localize memory pressure, improve cache behavior, and constrain the scope of garbage collection operations.
VGC further introduces a checkpoint-based bitfield tracking mechanism that encodes object state information compactly, enabling constant-time state evaluation using basic bitwise logic operations (e.g., AND, OR, and NOT). By replacing recursive heap traversal with checkpoint-based evaluation, the approach seeks to reduce garbage collection overhead and improve determinism. Additionally, a lightweight Yield Memory execution path is introduced to handle ephemeral operations—such as simple variable assignments and arithmetic expressions—without engaging full garbage collection logic.
This paper presents the architectural design of VGC, its zoning and checkpoint mechanisms, and its integration with a partition-aware execution model. Experimental observations are provided to illustrate the behavior of the proposed system under loop-intensive, recursive, and compute-heavy workloads. While the architecture represents an exploratory step rather than a fully realized replacement for Python’s existing memory management, the results suggest that zone-based allocation and partitioned execution provide a promising foundation for improving scalability and memory predictability in high-level language runtimes.
@@ 2 System Architecture
The Virtual Garbage Collector (VGC) is organized as a layered, zone-oriented architecture designed to promote deterministic memory behavior and controlled execution scalability. Rather than modifying Python’s existing memory manager directly, VGC is structured as an auxiliary architectural framework that explores alternative object lifecycle management strategies alongside partition-aware execution. At a high level, VGC replaces recursive heap traversal with a compact checkpoint-based bitfield representation for object state tracking. Object liveness and eligibility for reclamation are evaluated using constant-time bitwise operations, reducing reliance on per-object reference inspection during collection phases. This architectural choice aims to constrain garbage collection scope and improve predictability, particularly in loop-intensive and partitioned workloads. The system architecture separates execution partitioning from memory lifecycle control, allowing garbage collection decisions to be localized within predefined zones rather than applied globally across the heap. This separation provides a conceptual foundation for integrating memory zoning, checkpoint evaluation, and partition-aware execution without imposing changes on Python’s core interpreter semantics.
Detailed mechanisms and operational behavior of the zoning model, checkpoint logic, and execution integration are described in subsequent sections.
@@ 2.1 Overview of Architectural Layers
The Virtual Garbage Collector (VGC) is organized into four conceptual layers, each addressing a distinct aspect of object lifecycle management and execution control. The layers are designed to remain loosely coupled, allowing memory management decisions to be reasoned about independently from execution scheduling.
@@
    Memory Zoning Layer: Allocated objects are classified into one of three predefined memory zones—Red, Green, or Blue—based on observed access behavior and expected lifecycle characteristics. The zoning mechanism is intended to localize memory pressure and constrain garbage collection activity by grouping objects with similar usage patterns. Each zone maintains independent allocation metadata, reducing interference between workloads with differing memory access profiles.
@@
    Checkpoint Bitfield Layer: Instead of maintaining per-object reference counters, VGC employs a centralized checkpoint-based representation in which object states are encoded using compact bitfields. These bitfields represent coarse-grained lifecycle states (e.g., active, inactive, or eligible for reclamation). Object state evaluation is performed using constant-time bitwise logic operations such as AND, OR, and Logical NOT. This approach avoids recursive heap traversal and supports predictable state evaluation.
@@
    Yield Memory Layer: The Yield Memory layer provides a lightweight execution path for ephemeral operations, including temporary variable assignments and short-lived intermediate values. Objects handled within this layer may bypass full garbage collection tracking when their lifetime is limited to a single execution scope. Persistent values may optionally be registered with the checkpoint system once execution completes.
@@
    Control and Execution Layer: This layer coordinates execution partitioning and the interaction between memory zones, checkpoint evaluation, and yield memory handling. It encapsulates scheduling decisions and partition-aware execution control while remaining independent of low-level memory reclamation logic. The architecture distinguishes between runtime-oriented management (Active VGC) and compile-time allocation guidance (Passive VGC), with the latter treated as an intended optimization pathway rather than a fully realized component.
@@ 3.2 Yield Memory Execution Path
For short-lived and ephemeral operations, such as simple variable assignments,
x = 85
the standard garbage collection workflow may be intentionally avoided. The Yield Memory execution path provides a lightweight allocation context in which such operations can be evaluated without engaging full zoning or checkpoint evaluation mechanisms.
Within this execution path, temporary values are handled in a constrained scope and may be released immediately after execution completes. By limiting interaction with the broader memory management pipeline, Yield Memory reduces allocation overhead and avoids unnecessary garbage collection activity for trivial workloads.
When an object produced within the Yield Memory path must persist beyond its immediate execution scope, the system notifies the checkpoint mechanism, allowing the object to be registered and classified according to the active zoning policy. This transition enables ephemeral execution to coexist safely with longer-lived object management while preserving predictable memory behavior.
@@ 3.3 Three-Bit Checkpoint Architecture
The checkpoint layer constitutes the central mechanism for object state tracking within the Virtual Garbage Collector (VGC). In contrast to conventional reference counting schemes that maintain an integer counter for each object, VGC represents object state using a compact, fixed-width bitfield. This representation is intended to encode coarse-grained lifecycle information while enabling efficient state evaluation.
The checkpoint model uses a small number of bits to represent object status across evaluation cycles, allowing lifecycle transitions to be expressed explicitly without relying on recursive reference updates or heap traversal. Object state transitions are evaluated using constant-time bitwise logic operations, supporting predictable and bounded checkpoint processing.
By decoupling object liveness evaluation from per-object counters, the checkpoint architecture provides a structured alternative for exploring lifecycle-aware memory management. Detailed state semantics and transition rules are described in subsequent sections.
@@ 3.4 Object Feature Profiling and Runtime Metrics
  • $\rho(o)$Referential Fan-Out: The number of distinct inbound references points to the object, representing its connectivity within the program’s object graph.
@@ 3.10 Parallelism and Partitioning
VGC integrates a partitioning model that distributes workloads across CPU cores, with zones operating independently. Checkpoint bitfield updates are parallelized, avoiding locking overhead typical in Python’s GIL. Because the checkpoint system is based on 3-bit state logic, partitioning can use bitmask operations to batch-evaluate objects across zones concurrently. This allows multiple threads or interpreters to process memory in parallel without race conditions or serialized refcount updates. In summary, VGC’s working three-bit checkpoint state encoding, zone-based allocation, logic-gate-driven decision making, lambda-optimized yield execution, and parallel execution by partitioning workloads. This results in a deterministic, low-latency, and highly scalable alternative to Python’s GIL-based garbage collector.
@@ 4 Lambda and Three-Bit Checkpoint Mapping
5 Logic Gate–Based Checkpoint Architecture
The Logic Gate–Based Checkpoint Architecture[5], forms the computational core of the VGC checkpoint system. Rather than relying on iterative reference tracking or recursive object scanning, VGC encodes each object’s state within a compact 3-bit field, which is processed directly through gate-level logic operations. This design enables constant-time evaluation of object lifecycles and eliminates the overhead associated with conditional branching or traversal-based garbage collection.
@@ 5.1 Checkpoint Logic Integration
Each allocated object is associated with a 3-bit checkpoint code, represented as:
@@
Each bit encodes a specific state flag used to determine object liveness and readiness for reclamation. Rather than relying on traditional numerical reference counting, the proposed system groups checkpoint bitfields into logical blocks and evaluates them using bitwise logic operations. This design enables efficient, low-level state inspection across multiple objects simultaneously.
Object states are interpreted through combinations of logic gates that operate directly on checkpoint bits, allowing rapid binary-level evaluation of allocation status, zone permissions, and lifecycle transitions. The following logic gates are used within the checkpoint evaluation framework:
  • AND gate ($A\land B$) — Ensures that an object is considered live only when both its current state and corresponding zone permission bits are active. This gate enforces strict keep-alive conditions.
  • OR gate ($A\lor B$) — Preserves objects that remain associated with higher-priority zones (e.g., Green or Blue) even if their primary activation condition becomes inactive.
  • NOT gate ($\lnot A$) — Inverts inactive state bits to identify idle or expired objects that are eligible for reclamation.
  • XOR gate ($A\oplus B$) — Detects state transitions by identifying differences between successive evaluation epochs for the same object.
  • XNOR gate ($\overline{A\oplus B}$) — Identifies stability across multiple evaluation cycles, preventing premature reclamation of persistent objects.
  • NAND and NOR gates — Used in composite checkpoint circuits to support conditional sweeping and deferred reclamation. These gates are applied to objects marked with specific checkpoint patterns, such as $101$ (Deferred) and $110$ (Marked).
@@ 6.2 Core Partitioning Principle
where $P$ denotes the partition ratio representing the load distribution per core. If $T>C$, VGC dynamically clusters multiple Threads under a single core using an event-driven scheduler; if $T\leq C$, interpreters are mapped in a one-to-one configuration to maximize concurrency.
@@ 6.3 Zone-Based Partition Allocation
where $p$ represents a partition index. Each partition executes independently under a dedicated thread or lightweight core instance. The checkpoint and yield layers synchronize periodically to maintain zone integrity without introducing global locking mechanisms.
@@ 6.4 Parallel Execution Interpreter Synchronization
where $C_{T}$ denotes the computed checkpoint state for Thread $T$. Once local evaluations complete, the results are merged via a Sub-Allocator Aggregator that reorders and combines results from multiple threads and cores into a coherent memory snapshot.
@@ 6.6 Multi-Layer Parallel Model
  • Inter-Zone Parallelism: Red, Green, and Blue zones execute concurrently under separate threads or cores, merging results through the checkpoint synchronization layer.
@@ 6.8 Advantages of the Partition Model
6.9 Prototype Status and Research Scope
This work represents an initial-stage experimental investigation into partition-aware execution and zone-based memory management at the runtime-architecture level. The proposed PPE and VGC designs are evaluated as architectural prototypes intended to explore feasibility, behavioral properties, and design trade-offs, rather than as a production-ready replacement for existing Python runtimes.
The implementation focuses on validating core principles—such as deterministic partitioning, checkpoint-based lifecycle control, and strict space boundedness—under controlled workloads. As a result, several aspects commonly found in mature language runtimes, including full CPython integration, comprehensive language semantics, advanced optimizations, and ecosystem-level compatibility, are intentionally out of scope.
The goal of this work is to establish a clear architectural foundation and empirical baseline upon which future refinements, integrations, and optimizations can be built. Consequently, the results should be interpreted as evidence of architectural viability rather than as definitive performance claims for real-world Python applications.
@@ 8.7 Single core : No Partition 40K logical steps (deep recursion)
Table 2.7: Single-core PPE Deep Recursion Benchmark (40000 logical steps, recursion depth per chunk: 4000 steps)
@@ 8.8 Single core : No Partition 80K logical steps (deep recursion)
Table 2.8: Single-core PPE Deep Recursion Benchmark (80000 logical steps, recursion depth per chunk: 8000 steps)
@@ 8.9 Single core : No Partition 160K logical steps (deep recursion)
Table 2.9: Single-core PPE Deep Recursion Benchmark (160000 logical steps, recursion depth per chunk: 16000 steps)
@@ 8.19 Dual core : 2 Partitions 100K logical steps (deep recursion)
Table 2.19: Dual-core PPE Deep Recursion Benchmark (100000 logical steps, recursion depth per chunk: 10000 steps)
@@ 8.20 Dual core : 2 Partitions 200K logical steps (deep recursion)
Table 2.20: Dual-core PPE Deep Recursion Benchmark (200000 logical steps, recursion depth per chunk: 20000 steps)
@@ 8.21 Dual core : 2 Partitions 400K logical steps (deep recursion)
Table 2.21: Dual-core PPE Deep Recursion Benchmark (400000 logical steps, recursion depth per chunk: 40000 steps)